OpenStudy (anonymous):

MEDAL AND FAN TO BEST ANSWER: After tossing a number cube once find the probability of landing on a 3 or higher. P(3 or higher) = A. 2//3 B. 1/6 C. 1/2 D. 1/3

OpenStudy (anonymous):

So you want to obtain 3,4,5,6 out of 1,2,3,4,5,6

OpenStudy (anonymous):

4 out of 6 is like 2 out of 3

OpenStudy (anonymous):

So A is your answer

OpenStudy (anonymous):

ur ans is A. P(3 or higher)= 4/6 =2/3

OpenStudy (anonymous):

it is a.